THE APPLICATION OF A VUV-FT SPECTROMETER AND SYNCHROTRON RADIATION SOURCE TO MEASUREMENTS OF: I. PREDISSOCIATED LINE WIDTHS OF THE SCHUMANN-RUNGE BANDS OF O_{2}

\underline{K. YOSHINO}, J. R. ESMOND, W. H. PARKINSON, Harvard Smithsonian Center for Astrophysics, Cambridge, MA 02138, U.S.A.; A. P. THORNE, J. E. MURRAY, G. COX, R. C. M. LEARNER, Imperial College of Science, Technology, and Medicine, London, U.K.; K. ITO, T. MATSUI, Photon Factory, KEK, Tsukuba, Ibaraki 305, Japan; T. IMAJOU, Kyushu University, Fukuoka 812, Japan; A. S. -C. CHEUNG and K. -S. LEUNG, University of Hong Kong, Hong Kong.

The Imperial College VUV-FT spectrometer has been equipped with optically contacted, beam splitters made from single crystals of MgF_{2} and the short wavelength performance has been demonstrated down to ∼139 nm. To make ultrahigh resolution VUV photoabsorption cross section measurements with the VUV-FTS require a pure continuum source below 190 nm and the best choice: is synchrotron radiation from a storage ring facility. Moreover a suitable zero-dispersion predisperser is available on beam line 12-B of the synchrotron radiation source at the Photon Factory. We therefore moved the IC VUV FT spectrometer from Imperial College, London to the Photon Factory, Japan to exploit the bandwidthlimited synchrotron radiation as a background source for FT absorption spectroscopy. With resolution of 0.06 or 0.12 cm^{−1}, we extended the linewidth measurements of the Schumann-Runge (S-R) bands with v′ ≥ 12 in the wavelength range 181 to 175 nm. The observed line widths consist of the instrumental, Doppler, and predissociation widths. The line profiles resulting from the convolution of the first two of these contributions are very closely Gaussian with widths of 0.12 and 0.15 cm^{−1} for spectra recorded with resolutions of 0.06 and 0.12 cm^{−1}, respectively. The observed linewidths of the S-R bands with v′ ≥ 12 will be presented. This work is partly supported by NSF Division of Atmospheric Sciences grant ATM-94-22854 to Harvard College Observatory, and by NASA Upper Atmospheric Research Program under Grant No. NAG5-484 to the Smithsonian Astrophysical Observatory. We also acknowledge the support of the UK Science and Engineering Research Council and the Paul Instrument Fund of the Royal Society for the development of the UV-FT spectrometer.